# On Resolution of $\ell_{1}$-norm minimization via a Two-metric Adaptive Projection Method

### v1

@@ Abstract
The two-metric projection method is a simple yet elegant algorithm proposed by Bertsekas to address bound/box-constrained optimization problems. The algorithm’s low per-iteration cost and potential for using Hessian information make it a favorable computation method for this problem class. Inspired by this algorithm, we propose a two-metric adaptive projection method for solving the $\ell_{1}$-norm regularization problem that inherits these advantages. We demonstrate that the method is theoretically sound - it has global convergence. Furthermore, it is capable of manifold identification and has superlinear convergence rate under the error bound condition and strict complementarity. Therefore, given sparsity in the solution, the method enjoys superfast convergence in iteration while maintaining scalability, making it desirable for large-scale problems. We also equip the algorithm with competitive complexity to solve nonconvex problems. Numerical experiments are conducted to illustrate the advantages of this algorithm implied by the theory compared to other competitive methods, especially in large-scale scenarios. In contrast to the original two-metric projection method, our algorithm directly solves the $\ell_{1}$-norm minimization problem without resorting to the intermediate reformulation as a bound-constrained problem, so it circumvents the issue of numerical instability.
@@ 1 Introduction
The $\ell_{1}$-norm regularization is widely used in machine learning and statistics to promote sparsity in the solution. Many algorithms have been proposed to solve the $\ell_{1}$-norm regularization problem, such as the proximal gradient method [6], the accelerated proximal gradient method [18], the alternating direction method of multipliers (ADMM) [9], and the coordinate descent method [24]. However, these methods may converge slowly near the optimal point. Second-order methods are more desirable when seeking highly accurate solution. When $f$ is a twice continuously differentiable function, it is natural to expect algorithms to have faster convergence rates by exploiting the Hessian information of $f$ at each iterate $x_{k}$. Some successful algorithms have been proposed to solve the general composite optimization problem in the form of Eq. 1, such as the successive quadratic approximation (SQA) methods [30, 15], the semi-smooth Newton methods [14, 28] and the manifold optimization approach [3] based on the notion of partial smoothness [16].
In this work, we propose a two-metric adaptive projection method to address Eq. 1. The original two-metric projection method is proposed by Bertsekas [7, 12] to solve the bound constrained optimization problem:
@@
Two-metric projection method is a simple yet elegant algorithm which can be descriced as follows:
@@
where $D_{k}$ is a positive definite matrix, and $P(\cdot)$ denotes the Euclidean projection onto the nonnegative orthant. Note that $P(\cdot)$ has a simple closed form: given any vector $v\in\mathbb{R}^{n}$, $i$th component of $P(v)$ is $\max\{v^{i},0\}$. Therefore, the method has a low per-iteration cost if $D_{k}$ is provided. When $D_{k}$ is constructed using the Hessian information, the algorithm potentially converges fast. However, Bertsekas pointed out an issue regarding Eq. 3: starting from the current iterate $x_{k}$, the objective function value may not decrease for any small stepsize $\alpha_{k}>0$ even when $D_{k}$ is chosen as inverse of the Hessian. This means that a naive projected Newton step may fail. To solve this issue, Bertsekas proposed to divide the index set into two parts in each iteration:
@@
To avoid these issues, we propose a novel two-metric adaptive projection algorithm to solve Eq. 1. The algorithm is inspired by the two-metric projection method but directly solves Eq. 1 without resorting to the intermediate reformulation as a bound-constrained problem, so it circumvents the issue of numerical instability. Moreover, we do not require the Hessian to be uniformly positive definite nor solving the linear system exactly. Such design enables efficient and less stringent implementation to solve Eq. 1. Despite these improvements in numerical implementation, the algorithm preserves the identification of active set (optimal manifold) - it only needs to solve a low-dimensional linear system (Newton-equation) due to the sparsity structure of the optimal point. After manifold identification, it converges to the optimal solution in a superlinear rate. These convergence properties are shown using the strict complementarity and the error-bound condition that is weaker than the second-order sufficient condition and holds for a wide range of objective functions [32]. Our method is also favorable among the existing competitive second-order frameworks to solve Eq. 1. In particular, in each iteration we only need to solve a linear system of potentially low dimension inexactly, in sharp contrast to the proximal Newton methods [30, 15] which need to solve another $\ell_{1}$-norm regularized problem in full dimension at each step. Another method considered in [3] alternates between proximal gradient and Newton’s method to solve Eq. 1, aiming to use the former to identify the manifold and the latter to accelerate the convergence. The alternating design is necessary since they cannot know when a manifold is successfully identified. In contrast, our algorithm is able to identify the optimal manifold and reduce to a Newton’s method on the manifold automatically. These advantages are further verified via numerical experiments. We also explore the potential of this method to solve smooth nonconvex problems when the scaling matrix does not necessarily depend on Hessian. A competitve complexity close to the optimal $\mathcal{O}(\epsilon^{-2})$ is achieved after slight modification.
@@
    We provide a comprehensive theoretical analysis of the proposed algorithm, including global convergence, manifold identification and local superlinear convergence rates under the error bound condition and strict complementarity.
@@
  4. 4.
    We equip the algorithm with complexity guarantees close to the optimal when $f$ is smooth nonconvex.
The remainder of this paper is organized as follows: In Section 2 we clarify the basic assumptions and review the relevant background knowledge. In Section 3 the proposed algorithm is formally introduced in detail. Section 4 presents the theoretical analysis of the algorithm’s convergence properties. In Section 5 we discuss the experimental results. Discussion about the relationship between our algorithm and two-metric projection is included in Section 6, and we also illustrate in detail why we can not directly use the two-metric projection to solve the equivalent bound-constrained formulation of problem (1). Finally, in Section 7 we conclude the paper and discuss future work.
@@ 2.1 Nonsmooth optimization
If $h$ is a proper lower semicontinuous convex function, then $\operatorname{prox}_{h}(x)$ uniquely attained (c.f. [4]).
Relation between proximity operator and optimality conditions
If $x^{*}$ satisfies the first-order necessary conditions Eq. 5, we say that $x^{*}$ is a critical (stationary) point of Eq. 1. In particular, $x^{*}$ is critical iff
(6) $x^{*}=\operatorname{prox}_{\alpha h}(x^{*}-\alpha\nabla f(x^{*}))$
for any $\alpha>0$.
@@ 2.2 Riemannian optimization
For any $x\in\mathbb{R}^{n}$, define
$\mathcal{A}(x)\triangleq\left\{i\mid x^{i}=0\right\}.$
@@ 2.3 Basic assumptions
The following assumption will be used throughout the paper.
@@
$\psi=f+h$ is also lower semi-continuous therefore every sublevel set $\mathcal{L}_{\psi}\left(x_{0}\right)\triangleq\{x\mid\psi(x)\leq\alpha\}$ is closed thus compact. Therefore, $X^{*}$ is also compact.
@@
$x$ is a critical point that satisfies the strict complementarity condition if
$0\in\operatorname{relint}(\nabla f(x)+\partial h(x)).$
@@
$c_{1}\|x-\operatorname{prox}_{h}(x-\nabla f(x))\|\geq\operatorname{dist}\left(x,X^{*}\right)$
@@
For an optimal point $x^{*}\in X^{*}$, there exists $b_{1}>0$ such that for all $x\in\mathcal{M}_{*}\cap B\left(x^{*},b_{1}\right)$
@@
We illustrate the relationship between 2.8 and 2.9 via the following Theorem.
@@
Since $x^{*}$ satisfies SC, there exists a neighborhood $B\left(x^{*},2b_{1}\right)$ (w.l.o.g we suppose $b_{1}$ is same as 2.8) such that for all $x\in B\left(x^{*},2b_{1}\right)\cap X^{*}\subseteq\mathcal{M}_{*}$. Define $\Pi_{X^{*}}(x)\triangleq\operatorname{argmin}_{y\in X^{*}}\|y-x\|$, and let $\tilde{x}=\Pi_{X^{*}}(x)$. Then we have for all $x\in B\left(x^{*},b_{1}\right)$
$\|x-\tilde{x}\|=\operatorname{dist}\left(x,X^{*}\right)\leq\|x-x^{*}\|<b_{1}$
@@
$\|x-\tilde{x}\|=\operatorname{dist}\left(x,X^{*}\right)\leq\operatorname{dist}\left(x,X^{*}\cap\mathcal{M}_{*}\right)\leq\|x-\tilde{x}\|$
which means that $\operatorname{dist}\left(x,X^{*}\right)=\operatorname{dist}\left(x,X^{*}\cap\mathcal{M}_{*}\right)$. If $x\in\mathcal{M}_{*}$, then with a small neighborhood of $x^{*}$, we have
$\|x-\operatorname{prox}_{h}(x-\nabla f(x))\|=\|\operatorname{grad}_{\mathcal{M}_{*}}\psi(x)\|$
by SC and continuity (w.l.o.g. suppose this neighborhood is also $B\left(x^{*},b_{1}\right)$).
@@ 3 Two-metric Adaptive Projection Algorithm
Now we describe the two-metric adaptive projection algorithm for finding the critical point $x^{*}$ of Eq. 1. Analogously, we partition the index set into $I_{k}^{+}$ and $I_{k}^{-}$, and scale the gradient component in these two parts differently. In addition, the $I_{k}^{-}$ part needs to be further split to deal with the singularity of ||$\cdot$|| ⋅ | at 00. Definition of the projection operator depends on the current iterate, hence the name “adative projection”. Specifically, given $x_{k}\in\mathbb{R}^{n}$, accuracy level $\epsilon>0$, denote
@@
Let $\bar{g}_{k}$, $\bar{p}_{k}$ and $\bar{\omega}_{k,\epsilon}$ be the subvectors of $g_{k}$, $p_{k}$ and $\omega_{k,\epsilon}$ w.r.t index set $I_{k}^{-}$. Let $x_{k}^{+}$, $x_{k}^{+}(t)$, $g_{k}^{+}$ be the subvector of $x_{k}$, $x_{k}(t)$, $g_{k}$ w.r.t index set $I_{k}^{+}$, respectively. Define $h_{k}^{+}(y)\triangleq\gamma{\bf 1}^{T}|y|$, where $y\in\mathbb{R}^{|I_{k}^{+}|}$ and we apply ||$\cdot$|| ⋅ | to $y$ component-wisely.
@@
     if $\psi\left(x_{k}\right)-\psi\left(x_{k}\left(t_{k}\right)\right)<\sigma t_{k}(1-\tau)\mu_{k}\|\bar{p}_{k}\|^{2}+\sigma t_{k}\|G_{t_{k}}^{k}(x_{k}^{+})\|^{2}$ then
@@ 4 Convergence property
$I_{k}^{+}$ =$=\left\{1,2,\ldots,n_{1}\right\},$
$I_{k}^{-+}$ =$=\left\{n_{1}+1,\ldots,n_{2}\right\},$
$I_{k}^{--}$ =$=\left\{n_{2}+1,\ldots,n\right\}.$
@@
$\begin{pmatrix}G^{k}_{t}(x_{k}^{+})-g_{k}^{+}\\ \gamma\\ \vdots\\ \gamma\\ -\gamma\\ \vdots\\ -\gamma\end{pmatrix}\in\partial h(x_{k}(t))$
@@
(13) $h(x_{k})\geq h(x_{k}(t))+\left(G^{k}_{t}(x_{k}^{+})-g_{k}^{+}\right)^{T}tG_{t}^{k}(x_{k}^{+})+\gamma\sum_{i\in I_{k}^{-+}}(x_{k}^{i}-x_{k}^{i}(t))-\gamma\sum_{i\in I_{k}^{--}}(x_{k}^{i}-x_{k}^{i}(t))$
@@
$\leq\mu_{k}^{-1}(\|\bar{g}_{k}+\bar{\omega}_{k,\epsilon}\|+\|r_{k}\|)$
$\leq c^{-1}\|\bar{g}_{k}+\bar{\omega}_{k,\epsilon}\|^{1-\delta}+\tau\|\bar{p}_{k}\|$
@@
$\sum_{i\in I_{k}^{-}}\left((g_{k}^{i}+\omega_{k,\epsilon}^{i})(x_{k}^{i}(t)-x_{k}^{i})+\frac{L_{g}}{2}(x_{k}^{i}(t)-x_{k}^{i})^{2}\right)\leq\frac{L_{g}}{2}t^{2}\|\bar{p}_{k}\|^{2}-t(\bar{g}_{k}+\bar{\omega}_{k,\epsilon})^{T}\bar{p}_{k}$

### v2

@@ Abstract
The two-metric projection method is a simple yet elegant algorithm proposed by Bertsekas to address bound/box-constrained optimization problems. The algorithm’s low per-iteration cost and potential for using Hessian information make it a favorable computation method for this problem class. Inspired by this algorithm, we propose a two-metric adaptive projection method for solving the $\ell_{1}$-norm regularization problem that inherits these advantages. We demonstrate that the method is theoretically sound - it has global convergence. Furthermore, it is capable of manifold identification and has superlinear convergence rate under the error bound condition and strict complementarity. Therefore, given sparsity in the solution, the method enjoys superfast convergence in iteration while maintaining scalability, making it desirable for large-scale problems. Numerical experiments are conducted to illustrate the advantages of this algorithm implied by the theory compared to other competitive methods, especially in large-scale scenarios. In contrast to the original two-metric projection method, our algorithm directly solves the $\ell_{1}$-norm minimization problem without resorting to the intermediate reformulation as a bound-constrained problem, so it circumvents the issue of numerical instability.
@@ 1 Introduction
The $\ell_{1}$-norm regularization is widely used in machine learning and statistics to promote sparsity in the solution. Many algorithms have been proposed to solve the $\ell_{1}$-norm regularization problem, such as the proximal gradient method [6], the accelerated proximal gradient method [23], the alternating direction method of multipliers (ADMM) [9], and the coordinate descent method [29]. However, these methods may converge slowly near the optimal point. Second-order methods are more desirable when seeking highly accurate solution. When $f$ is a twice continuously differentiable function, it is natural to expect algorithms to have faster convergence rates by exploiting the Hessian information of $f$ at each iterate $x_{k}$. Some successful algorithms have been proposed to solve the general composite optimization problem in the form of Eq. 1, such as the successive quadratic approximation (SQA) methods [35, 16, 22], the semi-smooth Newton methods [15, 33, 20, 21] and the manifold optimization approach [3] based on the notion of partial smoothness and identifiablity [18, 17, 19].
In this work, we propose a two-metric adaptive projection method to address Eq. 1. The original two-metric projection method is proposed by Bertsekas and Gafni [7, 13] to solve the bound constrained optimization problem:
@@
Two-metric projection method is a simple yet elegant algorithm which can be described as follows:
@@
where $D_{k}$ is a positive definite matrix, and $P(\cdot)$ denotes the Euclidean projection onto the nonnegative orthant. Note that $P(\cdot)$ has a simple closed form: given any vector $v\in\mathbb{R}^{n}$, $i$th component of $P(v)$ is $\max\{v^{i},0\}$. Therefore, the method has a low per-iteration cost if $D_{k}$ is provided. When $D_{k}$ is constructed using the Hessian information, the algorithm potentially converges fast. However, there is an issue regarding Eq. 3: starting from the current iterate $x_{k}$, the objective function value may not decrease for any small stepsize $\alpha_{k}>0$ even when $D_{k}$ is chosen as inverse of the Hessian. This means that a naive projected Newton step may fail. To solve this issue, we need to divide the index set into two parts in each iteration:
@@
To avoid these issues, we propose a novel two-metric adaptive projection algorithm to solve Eq. 1. The algorithm is inspired by the two-metric projection method but directly solves Eq. 1 without resorting to the intermediate reformulation as a bound-constrained problem, so it circumvents the issue of numerical instability. Moreover, we do not require the Hessian to be uniformly positive definite nor solving the linear system exactly. Such design enables efficient and less stringent implementation to solve Eq. 1. Despite these improvements in numerical implementation, the algorithm preserves the identification of active set (optimal manifold) - it only needs to solve a low-dimensional linear system (Newton-equation) due to the sparsity structure of the optimal point. After manifold identification, it converges to the optimal solution in a superlinear rate. These convergence properties are shown using the strict complementarity and the error-bound condition that is weaker than the second-order sufficient condition and holds for a wide range of objective functions [37]. Our method is also favorable among the existing competitive second-order frameworks to solve Eq. 1. In particular, in each iteration we only need to solve a linear system of potentially low dimension inexactly, in sharp contrast to the proximal Newton methods [35, 16] which need to solve another $\ell_{1}$-norm regularized problem in full dimension at each step. Another method considered in [3] alternates between proximal gradient and Newton’s method to solve Eq. 1, aiming to use the former to identify the manifold and the latter to accelerate the convergence. The alternating design is necessary since they cannot know when a manifold is successfully identified. In contrast, our algorithm is able to identify the optimal manifold and reduce to a Newton’s method on the manifold automatically. These advantages are further verified via numerical experiments.
@@
    We provide a comprehensive theoretical analysis of the algorithm, including global convergence, manifold identification and local superlinear convergence rates under the error bound condition and strict complementarity.
@@
The remainder of this paper is organized as follows: In Section 2 we clarify the basic assumptions and review the relevant background knowledge. In Section 3 the proposed algorithm is formally introduced in detail. Section 4 presents the theoretical analysis. In Section 5 we discuss the experimental results. Discussion about two-metric projection to solve the equivalent bound-constrained formulation, relationship/comparison between this approach and ours are included in Section 6. Finally, in Section 7 we conclude the paper and discuss future work.
@@ 2.1 Nonsmooth optimization
If $x^{*}$ satisfies the first-order necessary conditions Eq. 5, we say that $x^{*}$ is a critical (stationary) point of Eq. 1.
@@
If $h$ is proper lower semicontinuous convex, then $\operatorname{prox}_{h}(x)$ uniquely attained (c.f. [4]).
Equivalent optimality conditions
$x^{*}$ is critical iff for any $\alpha>0$,
$x^{*}=\operatorname{prox}_{\alpha h}(x^{*}-\alpha\nabla f(x^{*})).$
@@ 2.2 Riemannian optimization
For any $x\in\mathbb{R}^{n}$, define $\mathcal{A}(x)\triangleq\left\{i\mid x^{i}=0\right\}.$
@@ 2.3 Basic assumptions
The following assumption will be used throughout.
@@
$\psi=f+h$ is lower semi-continuous therefore every sublevel set $\mathcal{L}_{\psi}\left(x_{0}\right)\triangleq\{x\mid\psi(x)\leq\alpha\}$ is closed thus compact. Therefore, $X^{*}$ is also compact.
@@
$x$ is a critical point that satisfies the strict complementarity condition if $0\in\operatorname{relint}(\nabla f(x)+\partial h(x)).$
@@
$c_{1}\|x-\operatorname{prox}_{h}(x-\nabla f(x))\|\geq\operatorname{dist}\left(x,X^{*}\right).$
@@
For an optimal point $x^{*}\in X^{*}$, there exists $b_{1}>0$ such that for all $x\in\mathcal{M}_{*}\cap B\left(x^{*},b_{1}\right)$,
@@
We illustrate the relationship between 2.8 and 2.9.
@@
Since $x^{*}$ satisfies SC, there exists a neighborhood $B\left(x^{*},2b_{1}\right)$ (w.l.o.g we suppose $b_{1}$ is same as 2.8) such that for all $x\in B\left(x^{*},2b_{1}\right)\cap X^{*}\subseteq\mathcal{M}_{*}$. Define $\Pi_{X^{*}}(x)\triangleq\operatorname{argmin}_{y\in X^{*}}\|y-x\|$, and let $\tilde{x}=\Pi_{X^{*}}(x)$. Then we have for all $x\in B\left(x^{*},b_{1}\right)$,
$\|x-\tilde{x}\|=\operatorname{dist}\left(x,X^{*}\right)\leq\|x-x^{*}\|<b_{1}.$
@@
$\|x-\tilde{x}\|=\operatorname{dist}\left(x,X^{*}\right)\leq\operatorname{dist}\left(x,X^{*}\cap\mathcal{M}_{*}\right)\leq\|x-\tilde{x}\|,$
which means that $\operatorname{dist}\left(x,X^{*}\right)=\operatorname{dist}\left(x,X^{*}\cap\mathcal{M}_{*}\right)$. If $x\in\mathcal{M}_{*}$ and is within a small neighborhood of $x^{*}$ (w.l.o.g. suppose this neighborhood is also $B\left(x^{*},b_{1}\right)$), we have
$i\in\mathcal{A}(x^{*})\Leftrightarrow x^{i}=0\Leftrightarrow|x^{i}-\nabla_{i}f(x)|<\gamma,$
$i\notin\mathcal{A}(x^{*})\Leftrightarrow x^{i}\neq 0\Leftrightarrow|x^{i}-\nabla_{i}f(x)|>\gamma,$
by SC and continuity. Therefore, in this neighborhood we have
$\|x-\operatorname{prox}_{h}(x-\nabla f(x))\|=\|\operatorname{grad}_{\mathcal{M}_{*}}\psi(x)\|,$
which concludes the proof.
@@ 3 Two-metric Adaptive Projection Algorithm
Now we describe the two-metric adaptive projection algorithm for finding the critical point $x^{*}$ of Eq. 1. Analogously, we partition the index set into $I_{k}^{+}$ and $I_{k}^{-}$, and scale the gradient component in these two parts differently. In addition, the $I_{k}^{-}$ part needs to be further split to deal with the singularity of ||$\cdot$|| ⋅ | at 00. Definition of the projection operator depends on the current iterate, hence the name “adaptive projection”. Specifically, given $x_{k}\in\mathbb{R}^{n}$, accuracy level $\epsilon>0$, denote
@@
Let $\bar{g}_{k}$, $\bar{p}_{k}$ and $\bar{\omega}_{k,\epsilon}$ be the subvectors of $g_{k}$, $p_{k}$ and $\omega_{k,\epsilon}$ w.r.t index set $I_{k}^{-}$. Let $x_{k}^{+}$, $x_{k}^{+}(t)$, $g_{k}^{+}$ be the subvector of $x_{k}$, $x_{k}(t)$, $g_{k}$ w.r.t index set $I_{k}^{+}$, respectively. Define $h_{k}^{+}(y)\triangleq\gamma\|y\|_{1}$, where $y\in\mathbb{R}^{|I_{k}^{+}|}$.
@@
     if $\psi\left(x_{k}\right)-\psi\left(x_{k}\left(t_{k}\right)\right)\geq\sigma t_{k}(1-\tau)\mu_{k}\|\bar{p}_{k}\|^{2}+\sigma t_{k}\|G_{t_{k}}^{k}(x_{k}^{+})\|^{2}$ then
@@ 4 Convergence property
$I_{k}^{+}=\left\{1,2,\ldots,n_{1}\right\},\;I_{k}^{-+}=\left\{n_{1}+1,\ldots,n_{2}\right\},\;I_{k}^{--}=\left\{n_{2}+1,\ldots,n\right\}.$
@@
$\left(G^{k}_{t}(x_{k}^{+})-g_{k}^{+},\gamma,\ldots,\gamma,-\gamma,\ldots,-\gamma\right)^{T}\in\partial h(x_{k}(t)).$
@@
(12) $h(x_{k})\geq h(x_{k}(t))+\left(G^{k}_{t}(x_{k}^{+})-g_{k}^{+}\right)^{T}tG_{t}^{k}(x_{k}^{+})+\gamma\sum_{i\in I_{k}^{-+}}(x_{k}^{i}-x_{k}^{i}(t))-\gamma\sum_{i\in I_{k}^{--}}(x_{k}^{i}-x_{k}^{i}(t)).$
@@
$\|\bar{p}_{k}\|$ $=\|(H_{k}+\mu_{k}I)^{-1}(\bar{g}_{k}+\bar{\omega}_{k,\epsilon}+r_{k})\|\leq\mu_{k}^{-1}(\|\bar{g}_{k}+\bar{\omega}_{k,\epsilon}\|+\|r_{k}\|)$
$\leq c^{-1}\|\bar{g}_{k}+\bar{\omega}_{k,\epsilon}\|^{1-\delta}+\tau\|\bar{p}_{k}\|,$
@@
$\sum_{i\in I_{k}^{-}}\left((g_{k}^{i}+\omega_{k,\epsilon}^{i})(x_{k}^{i}(t)-x_{k}^{i})+\frac{L_{g}}{2}(x_{k}^{i}(t)-x_{k}^{i})^{2}\right)\leq\frac{L_{g}}{2}t^{2}\|\bar{p}_{k}\|^{2}-t(\bar{g}_{k}+\bar{\omega}_{k,\epsilon})^{T}\bar{p}_{k}.$